You

A 1200 kg car moving at 12m/s applies it's break and stops . If the brakes are made of 5 kg of steel ( C = 450 JK ^-1 Kg^-1) each for 4 brakes ,by how much does the temperature of the metal rise ?



1
Expert's answer
2021-01-24T17:56:52-0500

First, let's write the kinetic energy lost due to braking:

"E = \frac{1}{2}mv^2"

Now, if we suppose that all this energy is converted to heat the breaks, we have:

"E = C\cdot m_{brake}\cdot4\cdot \Delta T"

By using the former expression:

"\frac{1}{2}m_{car}v^2 = C \cdot m_{brake}\cdot 4 \cdot \Delta T"

"\Delta T = \frac{m_{car}v^2}{8Cm_{brake}}=\frac{1200\cdot 144}{8\cdot450\cdot 5} = 9.6 ^{\circ}C (=9.6K)"